OpenStudy (anonymous):

3 cards are drawn from a standard deck of 52 cards. Find the probability of drawing either 3 red cards or 3 face cards.

OpenStudy (uri):

There are 12 face cards in a standard 52 card deck are 12/52=3/13 And 26 red cards -13 hearts and 13 diamonds. 26/52=1/2

OpenStudy (uri):

When we multiply both the probabilities it comes as 3/26 so it's 6/52.

OpenStudy (anonymous):

none of the answers i got

OpenStudy (amistre64):

P(red) = 1/2 P(face) = 4/13, or 3/13 .... depends on how you count an ace but some face cards are red .... so we have to adjust for that

OpenStudy (anonymous):

Probability (of drawing 3 red cards) 26/52 Probability (of drawing 3 face cards) 12/52 Probability (of drawing a face card and bring red) = 6/52 P (A or B) = P(A) + P(B) - P ( A and B) = 26/52 + 12/52 - 6/52 =

OpenStudy (uri):

I was almost there :D

OpenStudy (amistre64):

might want to adjust for "not replacing" maybe?

OpenStudy (anonymous):

Yes that is what I wanted to ask?

OpenStudy (anonymous):

Than you have to use a hyper geometric distribution...

OpenStudy (amistre64):

if we divvy it up int: r = red not face, R = red face, f = face not red .... we could brute it like this: rrr fff rrR rRr rRr rRR RrR RRr ffR fRf Rff fRR RfR RRf RRR

OpenStudy (anonymous):

P(red) = (26C1/52C1) x (25C1/51C1) x (24C1/50C1) = p(face) = (12C1/52C1) x(11C1/51C1) x (10C1/50C1) = p(face and red) = (6C1/52C1) x (5C1/51C1) x (4C1/50C1)) = Could you do it like this too @amistre64

OpenStudy (anonymous):

a.)84/137 b.)168/663 c.28/221 d.)120/1326

OpenStudy (anonymous):

Its C. just do what I did P(red) + P(face) - P(both) 2/17 + 11/1105 - 1/1105 = ? simplify..

OpenStudy (amistre64):

curious if my brute method pays out :) if we count aces as numbers: 20 red not face 6 red face 6 face not red ((20*19*18)+(6*5*4)+(6*5*4)+6(20*19*6)+6(6*5*6))/(52*51*50) = 14/85, thnx the wolf

OpenStudy (amistre64):

if we count aces as faces: 18 red not face 8 red face 8 face not red ((18*17*16)+2(8*7*6)+6(18*17*8)+6(8*7*8))/(52*51*50) = 956/5525 ... prolly not that way then

OpenStudy (amistre64):

slight error in my haste lol ... a counted rrR the same as RRr

OpenStudy (amistre64):

rrr (20*19*18) fff (6*5*4) rrR rRr 3(20*19*6) Rrr rRR RrR 3(20*6*5) RRr ffR fRf 3(6*5*6) Rff fRR RfR 3(6*5*6) RRf RRR (6*5*4) all over (52*51*50)